Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, mantovani, ruggi, casanueva, pinto, bersanetti, valentini, derossi, spinicelli, gosselin, gouaty - 18:20 Wednesday 01 December 2021 (54067) Print this report
ISC shift: SDB1 drift control, long lock

The goal of this shift was to test the effects of the drift control on SDB1, in conjunction with the alignment of the SR mirror.

In previous shifts was observed that tilting the SR on TY produced asymmetrical results, with TY+ being beneficial and TY- being almost immediately distruptive, This led us to believe that SDB1 could be misaligned wrt the ITF, and we decided to cooperate with DET people to engage its drift control loop, and confront the results with the misalignment of the SR.

Since the beginning of the shift locking the CITF was very difficult, and the lock acqusition attempts were distrupted by a 13-14 Hz oscillation at 160mW visible on most of the longitudinal DoFs, due to SRCL gain being too high (changed 0.8->0.65, tested and saved in the function). This considerably slowed down the tests in the first half of the shift.

To improve the stability of the PR alignment during the pre alignment, Ruggi swapped the optical lever used for the local control of the PR TY. This made the pre alignment process considerably easier, and we could proceed in the second half of the shift with the scheduled tests.

Closing SDB1 drift control on DC and SPRB drift control on 12MHz signals first, and 112MHz signals later, allowed us to achieve a lock that lasted 3h 40m (Fig.4). During this lock we:

  • closed the drift control of SDB1 on DC signal (immedaitely after locking the CITF)
  • changed the alignment of the SR from -1 to +1 urad wrt the initial point to investigate its effect
  • swapped the drift control of SPRB form 12MHz signal to 112MHz signal
  • towards the end, we tilted NE_TY in order to bring to 0 the dithering signals (this produced the power increase visible in the last minutes of the lock, in Fig.4)

As usual per the last few shifts, at carm offset zero we did not close any AA loop other than DIFFp and PR_X/Y, TX/Y, nor engaged SSFS.

In the figure is clearly visible an abrupt improvement of the quality of dark fringe, as well as the power on B7, B8 and B4 DC, or the shapes on the phase cameras, that became more symmetrical (Figs.2-3). This is due to the galvo loops on SDB2_B1p_QD2, that had opened without us realizing it, and closed again at that moment (Fig.1).

The lock was ultimately killed while changing the setpoint of NE_TY. The investigation of the unlock will be carried on in the next shifts, as well as the analysis of the data gathered during this lock.

We leave the ITF in DOWN to allow for the TCS actuators calibration.


Images attached to this report
Comments to this report:
tacca - 23:28 Wednesday 01 December 2021 (54070) Print this report
In the attached figures some data acquired by the phase cameras during the long lock (13:04 UTC and 16:42 UTC).
- Figure 1 shows the B4 images at the beginning of the lock -> 6 MHz and 56 MHz sidebands are characterized by an almost Gaussian shape;
- Figure 2 shows the B4 images just before the end of the lock -> 6 MHz and 56 MHz sidebands are characterized by an almost Gaussian shape.
The quality of the sidebands remained almost unchanged after more than 3 and a half hours of lock.
- Figure 3 shows some time data acquired by B1p and B4 phase cameras -> both on B1p and B4, after the usual large imbalance at the beginning of the lock, the 56 MHz USB and LSB are better balanced for most of the time of the lock, the imbalance increases after a couple of hours of lock; both on B1p and B4, the 6 MHz USB and LSB are almost always balanced for all the lock.
- Figure 4 shows a zoom of the same data at the beginning of the lock -> as already observed yesterday, on B1p USB and LSB cross a first time some seconds after the CARM null condition is reached and a second time after about 1 minute; while on B4 USB and LSB cross a first time at CARM null and a second time after about 1 minute.
Images attached to this comment
bonnand - 13:28 Thursday 02 December 2021 (54075) Print this report

Something worth noticing concerning the long lock of yesterday is that the cavity power was much lower than the previous attempts.

Figure shows the cavity power (B7/B8_DC) and the B4_12_MHz and B4_112MHz.

One can see that at the begining of the lock the power in the cavity follows the usual trend but after a few minutes dropped at 13:13 UTC and become noisier and then keep lowering up to the moment where the NE_Ty was moved (not visible in the plot).

I don't know what causes the cavity power to drop but it is not related to the opening of the galvo loop that happened 9 minutes later at around 13:22.

It is also noticeable that the B4 12MHz and 112MHz magnitude is increasing while the cavity power is decreasing but I cannot say if it is related.


Images attached to this comment
valentin - 17:10 Thursday 02 December 2021 (54080) Print this report

The cavity power was lower than usual most likely due to the arm dithering  and COMMp alignment loops being disabled, therefore their alignment not being optimal.

bonnand - 17:29 Thursday 02 December 2021 (54081) Print this report

I am putting here the plot of the error signals of the SPRB B4 QD2 quadrants during the long lock of yesterday.

  • Second row, are the horizontal signals for 12 MHz, 112Mhz and DC and for the bench position.
  • Third row is the same for vertical signals.
  • Fourth row are the signals of the PR and SR mirror in TX/TY.

The drift control of SPRB is on the 112 MHz signals starting from 13:23:21 UTC (it was 12 MHz before).

One can see that the horizontal DC channels is diverging other time whereas the vertical one is slowly getting closer to zero.

The 12 MHz channels are also drifting away from 0 in the same direction as the DC channels.

No evident correlation with the PR and SR mirrror channels can be seen.

Images attached to this comment
rocchi - 23:17 Tuesday 07 December 2021 (54144) Print this report

During the long lock of December 1st, it is possible to notice that, approximately after one hour from the beginning of the lock, the power in the sidebands started to increase (figure 1). At first glance, one would think that this is due to the slow decrease of the intracavity power (figure 2), which decreases the thermal lensing in the ITMs.

The trend of the sidebands power during this increase can be rather well fitted by an exponential function (see figure 3), which is the typical response to a step function load in case of thermal effects (the TM thermal inertia acts as a low pass filter). Thus, this behavior cannot be explained by the slow decrease of the intracavity power, which would lead to a linear increase of the sidebands. Another possibility could be a sudden change in the power of one of the TCS actuators, which is ruled out by figure 4.

As also suggested in VIR-1346A-21, a displacement of the ITF beam away from the ITM point absorbers cannot be excluded.

Images attached to this comment
Search Help